OpenStudy (anonymous):

How do you find the number of terms in an arithmetic sequence when only knowing the sum?

Parth (parthkohli):

The formula for sum is: \(\Large \color{MidnightBlue}{\Rightarrow {n \over 2}(a_1 + a_n) }\)

OpenStudy (ajprincess):

This is alternative one to find the sum. Sn=n/2 [2a + (n - 1)d]

OpenStudy (lgbasallote):

\[\large S_n = \frac{n(a_n + a_1)}{2}\] \[\large n = \frac {2S_n}{a_n+a_1}\]

Parth (parthkohli):

Duh!

OpenStudy (lgbasallote):

why didnt you think of that then @ParthKohli

Parth (parthkohli):

Lol I was making him work it out

OpenStudy (anonymous):

wait i'm interested in learning this ... could you please put in what each letter/variable represents.

OpenStudy (cwtan):

Sn= Sum of nth term a n= nth term

OpenStudy (anonymous):

so whats a represent?

OpenStudy (cwtan):

If the series 1,2,3,4,5 a1= First Term =1 a2=Second Term=2 an=nth Term If n =5 then an=5

OpenStudy (ajprincess):

d-common difference

OpenStudy (cwtan):

Oops I 4gt abt it

OpenStudy (ajprincess):

no problem.

OpenStudy (lgbasallote):

A1 means first term An means last term @aceace

OpenStudy (anonymous):

thanks... lol i am still to learn this stuff. is this topic in maths that difficult?

OpenStudy (cwtan):

Different for every person... But for me it's quite easy

OpenStudy (ajprincess):

Visit this link http://www.maths.com/sequences-and-series/arithmetic-sequences.htm. u will get some ideas on this, @aceace

OpenStudy (anonymous):

thank you guys ;)